OpenStudy (ttp):

Mark bought several containers of milk at the grocery store for $3 per gallon container. In this situation, what would the total cost of the milk be considered? domain range independent variable function

OpenStudy (ttp):

So can you help me @shamil98 ? :)

OpenStudy (shamil98):

f(x) = 3x the total cost is the function.

OpenStudy (ttp):

Thank you! :) Could you help me on some more? If not that's fine :)

OpenStudy (ttp):

Which of the following is a relationship between two variables in which the first variable determines exactly one second variable? function dependent variable relation ordered pair

OpenStudy (shamil98):

ordered pair

OpenStudy (ttp):

Which characteristic of a coordinate plane is false? A coordinate or point is written in the form ( x , y ). The quadrants are labeled one, two, three, and four. It consists of two quadrants and four axes. The two axes intersect at the origin.

OpenStudy (shamil98):

wait actually that one is a function... mb

OpenStudy (shamil98):

f(x) = y y = 3x let's say x = 1 y = 3(1) y = 3 so yeah its a function

OpenStudy (ttp):

It's Fine

OpenStudy (ttp):

Both of them?

OpenStudy (shamil98):

for the 3d one C is the false one. and yeah for both of them

OpenStudy (ttp):

Mhm k :) Jesse’s parents bought a new car. After owning it for two years, the car’s value was $20,085. After owning it for six years, Jesse’s parents sold it to him for $6,033. Make a graph, where x is the number of years they owned the car and y is the car’s value. Graph the car’s value at 2 years and 6 years. Determine the approximate purchase price of the car. $22,000 $24,000 $25,000 $27,000

OpenStudy (shamil98):

well for the first problem, the total cost = y. y = range. i'd go with range for that one , sorry, i'm a bit out of it today .

OpenStudy (ttp):

Ok :)

OpenStudy (shamil98):

after 2 years the car's price is 20,085 after 8 years the car's price is 6,033. around 14k difference in 6 years 14000 / 6 = 2333 so the yearly change is around 2 grand 2(2333) = 4666 20085 + 4666 = 24751 rounding up its about 25k as the original purchase price.

OpenStudy (ttp):

Mhm k :) I have still have some more can you help :)

OpenStudy (ttp):

A football team's kicker prepares to kick a field goal. The kicker stands at the origin of a coordinate plane. Facing the positive y-axis, the kicker takes 4 steps back, then 3 steps to the left. Which ordered pair best represents his position? four comma three three comma four negative three comma negative four negative four, negative three

OpenStudy (shamil98):

4 steps down, 3 steps left. x = the change horrizontally y = the change vertically. so..

OpenStudy (ttp):

So A?